1 litre mixture of CO and CO2 is taken. This is passed through a tube containing red hot charcoal. The volume now becomes 1.6 litre. The volumes are measured under the same conditions. Find the composition of mixture by volume.

Grade:Upto college level

4 Answers

Nandakumar UK
10 years ago

Suppose there is x litre CO and (1-x)L CO2. Here x litre (1-x)L CO2 reacts with C as C + CO2 --> 2CO. Then the final volume would be x + (1-x) + (1-x) =2 -x = 1.6; so x = 0.4L CO and 0.6L CO2

 

Regards :)

Priyanshu
6 years ago
Let the volume of CO=a litresAnd volume of CO2=b litres So,a+b=1And when it is passed through Charcoal tube, only CO2 will react,CO2 +C=2COb litres. 2b litresNow,a+2b=1.6And a+b= 1So,b= 0.6 litres and a is 0.4 litresAlso b=60%a=40%
Krish Gupta
3 years ago
 

Suppose there is x litre CO and (1-x)L CO2. Here x litre (1-x)L CO2 reacts with C as C + CO2 --> 2CO. Then the final volume would be x + (1-x) + (1-x) =2 -x = 1.6; so x = 0.4L CO and 0.6L CO

Thanks

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
Suppose there is x litre CO and (1-x)L CO2.
Here x litre (1-x)L CO2 reacts with C as C + CO2 →  2CO.
Then the final volume would be x + (1-x) + (1-x) =2 – x = 1.6;
so x = 0.4L CO and 0.6L CO2
 
Hope it helps.
Thanks and regards,
Kushagra